Circular No. 8386

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ds AND 2004dt Further to IAUC 8384, M. Moore and W. Li report the LOSS/KAIT

discovery of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004ds Aug. 11.47 2^{h}03^{m}56.^{s}51 −23^{o}18′50′′.6 16.8 2′′.0 W, 8′′.5 S
2004dt Aug. 11.48 2 02 12.77 − 0 05 51.5 16.1 7′′.0 E, 10′′.5 N

Additional approximate KAIT magnitudes: SN 2004ds in NGC 808, Jan. 29.15 UT, [19.0; Aug. 5.48, 16.8. SN 2004dt in NGC 799, Aug. 1.45, [18.0; 12.47, 15.9.

SUPERNOVA 2004dr IN ESO 479-G42 A. Gal-Yam and D. Fox, California Institute of Technology, report that

a red spectrum (range 550–780 nm) of SN 2004dr (cf. IAUC 8384), obtained by S. Kulkarni and Fox on Aug. 11.6 UT at the 10-m Keck I telescope (+ LRIS), shows a broad, well-developed P-Cyg Hα line and suggests that this is a type II-P supernova.

V1187 SCORPII R. W. Russell, D. K. Lynch, D. L. Kim, Aerospace Corporation; R. B.

Perry, Langley Research Center, NASA; M. L. Sitko, University of Cincinnati; T. R. Geballe, S. Fisher, G. Trancho, Gemini Observatory; T. H. Kerr, Joint Astronomy Centre; and W. Golisch, D. Griep, and P. Sears, University of Hawaii, report further infrared spectroscopy of V1187 Sco (cf. IAUC 8382). The H I line strengths measured with IRTF/BASS increased until Aug. 8 UT, when they peaked and then began a gradual decline. A higher-spectral-resolution 7–13-μm data set obtained with Gemini North (+ ‘Michelle’) on Aug. 11.25 revealed many H I lines but did not confirm the presence of the previously-reported, unidentified, infrared features at 8.6 and 11.2 μm (IAUC 8382). Those two features have now been shown to be due to H I 10-7 and 9-7, respectively. Additional BASS magnitudes on Aug. 11.18: L = 4.3; M = 3.9; narrowband N (10.5 μm) = 2.7.

SUPERNOVA 2004cs IN UGC 11001 CCD magnitudes of SN 2004cs (IAUC 8361) by A. Rajala, D. B. Fox,

and A. Gal-Yam on June 24.3 UT at the Palomar 1.5-m telescope: Johnson B = 18.16(7), V = 18.06(7), R = 17.87(8); Gunn g = 18.23(5).

2004 August 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green